# Can Semantic Methods Enhance Team Sports Tactics? A Methodology for Football with Broader Applications

### v1

@@ 6.3 Match Observations
Table 6: Observed team attributes for SSV Pachten across both match halves.
Attribute First Half Second Half $\Delta$
@@
Offensivkraft ($A_{1}$) Hoch 0.85 Hoch 0.85 0.00
Direkte vertikale Angriffe ($A_{4}$) Hoch 0.85 Mittel 0.50 0.35-0.35
Gegenangriff ($A_{4}$) Hoch 0.85 Hoch 0.85 0.00
Kompakte Defensive ($A_{2}$) Mittel 0.50 Niedrig 0.20 0.30-0.30
Restenergie ($A_{8}$) Mittel 0.50 Niedrig 0.20 0.30-0.30
Gegenpressing ($A_{5}$) Mittel 0.50 Mittel 0.50 0.00
@@ 6.4.2 Strategy Comparison
Table 7 presents the adapted semantic distances between the projected team vector and the subset of strategy templates relevant to the observable attribute space.
@@ 6.5 Retrospective Analysis
Attribute DSS Rec. Observed Alignment
Offensive Strength ($A_{1}$) 0.80 0.85 $\checkmark$
Defensive Strength ($A_{2}$) 0.50 0.20 $\times$
Transition Speed ($A_{4}$) 0.50 0.85 $\times$
High Press Capability ($A_{5}$) 0.40 0.50 $\checkmark$
Residual Energy ($A_{8}$) 0.60 0.20 $\times$
@@ 6.7 Implications for Framework Validation
Despite its limitations, this pilot study demonstrates several important capabilities that inform the broader research agenda:
  • Real-data compatibility: The DSS can ingest observational data from actual matches using a straightforward categorical-to-continuous mapping protocol.
  • Temporal dynamics: The framework successfully captures intra-match evolution (first half $\to$ second half), enabling phase-specific recommendations.
  • Diagnostic interpretability: The attribute-level analysis provides actionable insights (e.g., “energy reserves constrain high-intensity options”) that coaches can readily interpret.
  • Graceful degradation: Even with partial attribute coverage (5 of 14 dimensions), the DSS produces coherent recommendations, suggesting robustness to incomplete information.
@@
This pilot case study represents an essential step in the research trajectory: from theoretical formalization (Section 3) through prototype implementation (Section 4) and controlled experimentation (Section 5) to real-world application. While preliminary, the results demonstrate that the semantic-distance approach can accommodate observational data from actual matches while preserving interpretability and adaptability. The following Discussion (Section 7) synthesizes insights from both the simulated experiments and this pilot study, reflecting on limitations and charting directions for future development.
@@ 7 Discussion
The experimental evaluation demonstrates that the proposed semantic-distance Decision Support System achieves a high degree of internal coherence and provides tactically meaningful recommendations across heterogeneous match scenarios. The system exhibits both stability—particularly in balanced or high-energy contexts—and interpretability through its diagnostic visualizations. Nevertheless, beyond the pilot-specific constraints noted in Section 6.6, the DSS architecture itself presents broader limitations that constrain the current prototype’s applicability and operational readiness.
@@ 7.1 Methodological Limitations
Data quality and representativeness.
@@
Static opponent modelling.
@@
Linear distance assumptions.
@@
Absence of operational constraints.
@@
User-facing interpretability.
@@
8 Conclusion and Future Work
This work introduced a Decision Support System for context-aware football strategy selection, grounded in a semantic model that represents both teams and strategies as vectors in a shared 14-dimensional attribute space. The adjusted semantic-distance metric combines static team–strategy compatibility with dynamic contextual factors—match time, score state, residual energy, and opponent adaptability—controlled by explicit weighting functions. Validation through synthetic scenarios and a pilot study with real match data demonstrated that the DSS produces coherent recommendations, identifies the factors driving each decision, and degrades gracefully when only partial attribute coverage is available.
@@ 8.1 Summary of Contributions
    An adaptive distance metric that dynamically reweights attributes based on match context (energy, time pressure, opponent gaps), with explicit, reproducible formulas.
@@
    Diagnostic interpretability tools—radar charts, sensitivity analysis, robustness testing—that expose the reasoning behind recommendations.
@@
    Pilot validation with real match data, demonstrating applicability beyond synthetic scenarios.
@@ 8.2 Future Directions
Advanced data integration and modeling.
@@
Non-linear and hybrid metrics.
@@
Multi-objective optimization.
@@
Interactive coaching interface.
@@
Validation with professional data.
@@
Extension to other team sports.
@@
From strategy selection to strategy synthesis.
@@
Adversarial and security domains.

### v2

@@ 6.4 Match Observations
Table 16: Observed team attributes for SSV Pachten across both match halves. German source terms are shown with corresponding DSS identifiers per the mapping in Table 15. Categorical values (Hoch/Mittel/Niedrig) are converted to normalized scores via Equation (14).
Attribute (German $\to$ DSS ID) First Half Second Half $\Delta$
@@
Offensivkraft $\to$ $A_{1}$ Hoch 0.85 Hoch 0.85 0.00
Direkte vert. Angriffe $\to$ $A_{4}$ Hoch 0.85 Mittel 0.50 0.35-0.35
Gegenangriff $\to$ $A_{4}$ Hoch 0.85 Hoch 0.85 0.00
Kompakte Defensive $\to$ $A_{2}$ Mittel 0.50 Niedrig 0.20 0.30-0.30
Restenergie $\to$ $A_{8}$ Mittel 0.50 Niedrig 0.20 0.30-0.30
Gegenpressing $\to$ $A_{5}$ Mittel 0.50 Mittel 0.50 0.00
@@ 6.5.2 Strategy Comparison
Table 17 presents the adapted semantic distances between the projected team vector and the subset of strategy templates relevant to the observable attribute space. Figure 6 visually compares the adaptability of these strategies to the team’s current condition.
@@ 6.6.2 Counterfactual Consideration
6.7 Evaluation Results
This section reports outcomes against the evaluation endpoints and baselines specified in Section 6.1.
6.7.1 Endpoint 1: Processing Feasibility
Result: PASSED.
The DSS successfully processed the categorical observational data through the complete pipeline:
  • Categorical-to-numerical conversion executed without errors for all 6 observed attributes.
  • Reduced 5-dimensional team vector constructed (after aggregating $A_{4}$ from two sources).
  • Semantic distances computed for all 20 strategies in the library.
  • Ranked recommendation list generated with diagnostic output for top-3 strategies.
  • Halftime projection with fatigue adjustment produced valid results.
No degenerate outputs (e.g., tied rankings, infinite distances, missing values) were observed. Pipeline execution time was ¡50 ms, confirming suitability for real-time deployment.
6.7.2 Endpoint 2: Expert Agreement
Result: PASSED (with partial agreement on top-1).
Two independent expert reviewers with football coaching experience evaluated the DSS recommendation, with outcomes summarized in Table 18:
Table 18: Expert assessment of DSS recommendation for the pilot match.
Assessment Item Expert 1 Expert 2
Top-1 recommendation appropriate? Appropriate Partially Appropriate
Top-3 contains endorsed strategy? Yes (Build-up Play) Yes (Cautious Horizontal)
Would use DSS output in practice? Yes, with caveats Yes, as input to discussion
Qualitative feedback:
  • Expert 1: “Build-up Play is a sensible choice given the energy constraints. The diagnostic correctly identifies stamina as the limiting factor. I would have recommended the same.”
  • Expert 2: “Build-up Play is reasonable but perhaps too conservative for a team leading 2:1 at halftime. I might prefer Cautious Horizontal Play [ranked 3rd by DSS] to maintain some attacking threat. However, the DSS’s reasoning is sound and the top-3 set is useful.”
The success criterion (both experts rate top-1 as Appropriate or Partially Appropriate; both endorse at least one top-3 strategy) was met.
6.7.3 Endpoint 3: Tactical Alignment Analysis
Result: Divergence observed; descriptive analysis provided.
As reported in Table 19, the team’s second-half tactics diverged from the DSS recommendation on 3 of 5 observable dimensions. The team maintained high-intensity, transition-focused play despite declining energy reserves—a higher-risk approach than the DSS recommended.
Attribute DSS Rec. Observed Alignment
Offensive Strength ($A_{1}$) 0.80 0.85 $\checkmark$
Defensive Strength ($A_{2}$) 0.50 0.20 $\times$
Transition Speed ($A_{4}$) 0.50 0.85 $\times$
High Press Capability ($A_{5}$) 0.40 0.50 $\checkmark$
Residual Energy ($A_{8}$) 0.60 0.20 $\times$
Outcome association (descriptive only):
  • The team won the match (4:3), suggesting the high-risk approach succeeded in this instance.
  • However, the team conceded 2 second-half goals (vs. 1 in the first half), consistent with the DSS’s warning about defensive vulnerability under energy depletion.
  • Causal attribution is not possible from a single match.
6.7.4 Baseline Comparisons
Table 20 provides a baseline comparison of recommendation methods.
Table 20: Baseline comparison of recommendation methods.
Method Recommendation Expert 1 Expert 2
DSS (proposed) Build-up Play Appropriate Partially Appr.
Random baseline Offside Trap Inappropriate Inappropriate
Default strategy Build-up Play Appropriate Partially Appr.
Energy-only heuristic Positional Defense Partially Appr. Inappropriate
  • Random baseline: Selected “Offside Trap” (a high-risk defensive tactic requiring precise coordination)—deemed inappropriate by both experts for a fatigued youth team.
  • Default strategy baseline: Coincidentally matched the DSS recommendation (Build-up Play), achieving the same expert ratings. This highlights that Build-up Play is indeed a “safe” choice but also that the DSS does not always outperform simple defaults. The DSS’s value lies in (i) providing diagnostic reasoning and (ii) adapting to contexts where the default would be inappropriate.
  • Energy-only heuristic: Recommended “Positional Defense” based solely on low projected energy ($A_{8}=0.35<0.4$). Expert 1 rated this as “Partially Appropriate” (energy conservation is valid), but Expert 2 rated it “Inappropriate” (too passive for a team with strong offensive capability and a lead to protect through controlled possession rather than pure defense).
Key insight: The DSS matched the default baseline in this scenario but provided richer diagnostic output. The energy-only heuristic, while capturing one important factor, missed the interaction between offensive capability and energy management that the full DSS model captures.
6.7.5 Summary of Evaluation Outcomes
The pilot study achieved its primary and secondary endpoints, as summarized in Table 21. The DSS successfully processed real observational data and produced recommendations that independent experts judged appropriate or partially appropriate. Performance matched the default baseline but exceeded the random and energy-only baselines, with the added benefit of diagnostic interpretability.
Table 21: Summary of pilot study evaluation outcomes.
Endpoint Criterion Outcome
1. Processing Feasibility Pipeline executes without errors PASSED
2. Expert Agreement Both experts: Appr. or Part. Appr. PASSED
3. Tactical Alignment Descriptive comparison Divergence observed
Baseline: Random Expert agreement Failed (0/2)
Baseline: Default Expert agreement Matched DSS (2/2)
Baseline: Energy-only Expert agreement Partial (1/2)
@@ 6.9 Implications for Framework Validation
Despite its limitations, this pilot study provides preliminary evidence for several framework capabilities, though these observations should be interpreted cautiously given the single-match sample and partial attribute coverage:
  • Real-data compatibility: The DSS can ingest observational data from actual matches using a straightforward categorical-to-continuous mapping protocol (see Section 6.3), suggesting that the framework is not inherently limited to synthetic inputs.
  • Temporal dynamics: The framework captures intra-match evolution (first half $\to$ second half), enabling phase-specific recommendations. Whether this capability generalizes across match contexts remains to be established.
  • Diagnostic interpretability: The attribute-level analysis provides insights (e.g., “energy reserves constrain high-intensity options”) that appear actionable, though coach acceptance testing has not been conducted.
  • Graceful degradation: Even with partial attribute coverage (5 of 14 dimensions), the DSS produces coherent recommendations. This robustness to incomplete information is encouraging but requires systematic evaluation across varying degrees of data availability.
Important caveat: These observations demonstrate feasibility rather than validity. The pilot shows that the DSS can process real match data and produce interpretable outputs; it does not establish that these outputs would improve coaching decisions or match outcomes. Operational validity claims require prospective studies with systematic outcome tracking.
@@
This pilot case study represents one step in the research trajectory: from theoretical formalization (Section 3) through prototype implementation (Section 4) and controlled experimentation (Section 5) toward real-world application. The results establish that the semantic-distance approach can accommodate observational data from actual matches while preserving interpretability and adaptability. However, a single retrospective application cannot establish operational validity or predictive value. The following Discussion (Section 7) synthesizes insights from both the simulated experiments and this pilot study, explicitly delimiting the scope of current claims and charting the validation work required before stronger conclusions can be drawn.
@@ 7 Discussion
The experimental evaluation and pilot study provide evidence that the proposed semantic-distance Decision Support System achieves internal coherence within its design parameters and produces recommendations that align with tactical intuition across the tested scenarios. Following revisions to ensure method–implementation consistency (Sections 3.54.2), establishment of a systematic localization policy (Section 6.3), and provision of complete formal specifications (Appendix A), the analytical pipeline from input data to tactical recommendations is now fully auditable and reproducible.
However, auditability does not imply operational validity. The system exhibits stability in balanced or high-energy contexts and interpretability through its diagnostic visualizations, but these properties have been demonstrated primarily within synthetic evaluation frameworks. Beyond the pilot-specific constraints noted in Section 6.8, the DSS architecture itself presents broader limitations that constrain claims regarding real-world applicability and operational readiness.
@@ 7.1 Methodological Limitations
7.1.1 Data Quality and Representativeness
@@
7.1.2 Static Opponent Modelling
@@
7.1.3 Linear Distance Assumptions
@@
7.1.4 Absence of Operational Constraints
@@
7.1.5 User-Facing Interpretability
@@
8 Conclusions and Future Work
This work introduced a Decision Support System for context-aware football strategy selection, grounded in a semantic model that represents both teams and strategies as vectors in a shared 14-dimensional attribute space. The adjusted semantic-distance metric combines static team–strategy compatibility with dynamic contextual factors—match time, score state, residual energy, and opponent characteristics—controlled by explicit weighting functions documented in full (Appendix A).
Evaluation through synthetic scenarios demonstrated internal consistency: the DSS produces recommendations that align with tactical intuition, responds appropriately to contextual variations, and provides interpretable diagnostics. A pilot study with real match data established feasibility of processing observational inputs, though the single-match sample and partial attribute coverage preclude claims of generalizability. Critically, these results demonstrate that the system behaves as designed; whether DSS recommendations would improve actual coaching decisions or match outcomes remains an open empirical question requiring prospective validation.
@@ 8.1 Summary of Contributions
    An adaptive distance metric that dynamically reweights attributes based on match context (energy, time pressure, opponent gaps), with explicit, reproducible formulas (Section 3.6.2, Appendix A.2).
@@
    Diagnostic interpretability tools—radar charts, sensitivity analysis, robustness testing, ablation studies—that expose the reasoning behind recommendations and enable systematic evaluation.
@@
    A fully auditable pipeline with complete formal specifications, code availability, and localization protocols that support independent replication and verification.
  5. 5.
    Preliminary real-data application via a pilot study, demonstrating feasibility (though not yet validity) of processing observational match data.
@@ 8.2 Future Directions
8.2.1 Advanced Data Integration and Modeling
@@
8.2.2 Non-Linear and Hybrid Metrics
@@
8.2.3 Multi-Objective Optimization
@@ 8.2.4 Predictive Simulation
8.2.5 Interactive Coaching Interface
@@
8.2.6 Validation with Professional Data
@@
8.2.7 Extension to Other Team Sports
@@
8.2.8 From Strategy Selection to Strategy Synthesis
@@
8.2.9 Adversarial and Security Domains